OpenStudy (anonymous):

which of the following graphs represents the equation y= -2x + 3 ?

OpenStudy (anonymous):

@jim_thompson5910 @precal

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@JamesJ

OpenStudy (anonymous):

@whpalmer4 @Luis_Rivera

OpenStudy (anonymous):

@timo86m @OakTree

OpenStudy (anonymous):

B y=mx+b y=-2x + 3 the one that goes down 2 over 1. and y intercept at 3

OpenStudy (anonymous):

I got D? @Luis_Rivera @Hero

OpenStudy (anonymous):

graph B

OpenStudy (anonymous):

@OakTree

OpenStudy (anonymous):

okay ? @Luis_Rivera Thank you! @k.rajabhishek

OpenStudy (whpalmer4):

@needmathhelp7 Use the x and y-intercepts to decide which graph. y-intercept is where x = 0, so y = -2(0) + 3 = 3. Any graph that doesn't show the line crossing through the point (0,3) can be rejected. x-intercept is where y = 0, so 0 = -2x+3, 2x = 3, x = 1.5. Any graph that doesn't show the line crossing through the point (1.5,0) can be rejected. In combination, that leaves only graph B.